A student begins with 9.6 mL of a 0.28 M Cu(NO3)2 solution and performs the copper cycle lab. She recovers 0.070 grams of copper metal.

What is the theoretical yield of copper in grams?
Answer #1

volume , V = 9.6 mL

= 9.6*10^-3 L

use:

number of mol,

n = Molarity * Volume

= 0.28*9.6*10^-3

= 2.688*10^-3 mol

This is number of moles of Cu(NO3)2

one mole of Cu(NO3)2 contains 1 moles of Cu

use:

number of moles of Cu = 1 * number of moles of Cu(NO3)2

= 1 * 2.688*10^-3

= 2.688*10^-3

Molar mass of Cu = 63.55 g/mol

use:

mass of Cu,

m = number of mol * molar mass

= 2.688*10^-3 mol * 63.55 g/mol

= 0.1708 g

Answer: 0.171 g